OpenStudy (anonymous):

If 3x^2= 8y = 48, what is the value of x^2y

OpenStudy (anonymous):

if \(8y=48\) then \(y=48\div 8=6\)

OpenStudy (anonymous):

Because 8y=48, then y= ?? can you answer this simple question ?

OpenStudy (anonymous):

and if \(3x^2=48\) then \(x^2=48\div 3\)

OpenStudy (anonymous):

yes i can answer that SIMPLE question. its 6.

OpenStudy (anonymous):

making \(x^2=16\) and so \(x^2y=16\times 3\)

OpenStudy (anonymous):

3x^2=48 x^2=48/3=16 8y=48 y=48/8=6 x^2*y=16*6=?